# Adaptive Attention Distillation for Robust Few-Shot Segmentation under Environmental Perturbations

### v1

@@ IV-F Decoder and Loss Function
The final stage of the AAD framework is to synthesize the information gathered by the preceding modules into a precise segmentation map. As illustrated in Fig. 3(d), this is achieved by the decoder, which integrates the coarse localization from the Correlation Learner with the high-level semantic knowledge from the AAD learner.
The inputs to this stage are the multi-scale coarse query masks $M^{q^{\prime}}_{i}$ and the final set of discriminative class queries, $\hat{q}$. To effectively combine these, the class queries $\hat{q}$ are used to refine the coarse masks. Specifically, we perform an element-wise multiplication between $\hat{q}$ and each coarse mask $M^{q^{\prime}}_{i}$. This operation re-weights the mask features, amplifying regions that are consistent with the distilled class semantics and suppressing irrelevant areas. The resulting refined masks are then concatenated with the original coarse masks to form a rich, fused representation, $R_{q}$, as shown in Equation 4:
@@
where $\otimes$ denotes element-wise multiplication.
Finally, this fused representation $R_{q}$ is processed by the decoder module. The decoder is a simple yet effective architecture composed of several convolutional and upsampling layers that progressively merge the multi-scale features and restore the representation to the original image resolution, producing the final segmentation prediction. The entire network is trained end-to-end. We optimize the model parameters by minimizing the standard Binary Cross-Entropy (BCE) loss between the predicted segmentation map and the ground-truth query mask.
@@ V Experiments
ResNet-50 HSNet [ICCV2021] [32] 2.25 2.58 3.65 6.35 6.72 16.46 18.75 20.39 20.63
CyCTR [NIPS2021] [49] 0.88 0.89 0.83 7.98 9.24 3.01 13.05 13.56 11.37
PFENet  [TPAMI2022] [42] 0.44 0.41 0.39 9.06 9.06 9.14 11.60 11.06 11.13
@@
ResNet-101 CyCTR [NIPS2021] [49] 0.20 0.02 0.07 5.20 3.97 9.89 17.18 16.44 16.14
RepriNet [CVPR2021] [2] 1.31 4.13 1.47 6.70 5.54 3.04 12.40 11.53 9.86
DCAMA [ECCV2022] [39] 1.55 1.60 1.62 8.08 9.06 17.52 22.91 26.47 28.23
@@
ResNet-50 HSNet [ICCV2021] [32] 45.78 48.91 49.14 9.57 9.61 9.60 15.42 14.91 14.52
CyCTR [NIPS2021] [49] 46.18 47.76 44.25 9.06 9.15 9.00 12.93 12.88 12.81
PFENet  [TPAMI2022] [42] 32.89 33.43 32.74 9.06 9.06 9.14 13.48 13.73 13.81
@@
ResNet-101 CyCTR [NIPS2021] [49] 52.20 52.36 53.49 9.12 9.03 8.99 13.40 12.68 13.06
RepriNet [CVPR2021] [2] 46.21 49.84 49.74 9.88 11.05 12.36 16.41 23.73 30.97
DCAMA [ECCV2022] [39] 55.51 59.35 59.06 10.20 10.75 11.21 21.71 25.13 31.85
@@
TABLE III: Comparison with SOTA methods on 1-shot, 5-shot, and 20-shot setting on Lunar terrain (Astronomy), and City satellite (Geography) datasets. The ’Mean’ refers to the average results across all eight datasets in the FSHS benchmark. The numbers in Bold indicate the best performance.
@@
ResNet-50 HSNet [ICCV2021] [32] 3.84 4.01 4.15 9.61 10.08 10.25 14.98 14.65 16.05
CyCTR [NIPS2021] [49] 4.78 5.70 5.42 6.64 6.69 6.33 12.98 13.84 12.19
PFENet  [TPAMI2022] [42] 7.02 7.65 7.98 4.91 4.66 4.50 11.06 11.13 11.10
@@
ResNet-101 CyCTR [NIPS2021] [49] 4.94 5.64 6.28 5.40 5.06 4.69 13.46 13.15 14.08
RepriNet [CVPR2021] [2] 3.41 9.00 12.35 8.74 10.33 10.65 13.13 15.60 16.35
DCAMA [ECCV2022] [39] 4.91 7.62 8.23 9.48 10.00 11.25 16.79 20.03 21.12
@@ V-A Experiment Setup
Training and Testing Strategy. For pre-trained segmentation models (SAM [20]), we adhere to their inherent transfer learning training strategy by conducting end-to-end training on the pre-training dataset for the entire model. During the evaluation phase, we fine-tune the pre-trained model using support images, followed by generating segmentation predictions for query images. Regarding FSS models (HSNet [32], CyCTR [49], PFENet [42], DCAMA [39], DIaM [16], HDMNet [34], RepriNet [2], SCCAN [48],PFENet++ [30], HMNet [47], ABCDFSS [18], NTRENet++ [26] ), we adopt a meta-learning training strategy and subsequently evaluate the trained models using a meta-testing strategy. In each evaluation, we compute the average mean-IoU over $2$ runs [43], each with different random seeds. Additionally, each run comprises $1,000$ tasks for each dataset across the evaluation benchmark, maintaining consistency with the setting in  [24].
@@
Implementation Details. For a fair comparison, we employ Swin-transformer [27], ResNet-50 [17], and ResNet-101 [17] as feature extraction networks, all of which are initialized with the weights pre-trained on ILSVRC [36] and kept frozen during the training process, following the previous works [24, 2]. The input dimensions for support and query images are set at $384\times 384$. ResNet-50 and ResNet-101 feature maps have channel dimensions of 256, 512, 1024, and 2048, while Swin-transformer feature maps have dimensions of 192, 384, 768, and 1536. For ResNet-50-based ADD, the number of learnable initialization queries is set to 15, and the interaction with $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ dimensions of support and query features. In the case of the ResNet-101-based model, the number of queries is set to 20, with feature interactions at $\frac{1}{32}$ dimension Meanwhile, for the Swin-transformer-based model, 15 queries are used, with feature interactions at $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ dimensions. The decoder is configured with $2$ convolutional layers, and between each module, bilinear interpolation is applied to upsample the feature maps by a factor of $2$, resulting in a total of $2$ upsampling functions. These networks were implemented using PyTorch, with AdamW [29] as the optimizer, a learning rate of 1e41e-4, and a weight decay of $0.05$. During training, the batch size is set to $120$, and the training process ran on $8$ NVIDIA A800-SXM4-80GB GPUs in parallel, with subsequent evaluation on one GPU.
@@ V-B Comparison with SOTA Models
As in Tab. III, and III, extensive evaluations across the eight datasets of the ER-FSS benchmark demonstrate that our proposed AAD framework consistently and substantially outperforms existing SOTA FSS and pretrain-finetune models across all backbones and settings. When using a ResNet-50 backbone, AAD achieves a mean mIoU of 19.24%, 21.90%, and 23.84% in the 1-shot, 5-shot, and 2 0-shot settings, respectively. This represents a significant margin over the next best method, DCAMA, which scores 14.60%, 15.99%, and 17.09%. The performance gap is even more pronounced with a Transformer backbone, where AAD achieves a mean mIoU of 23.47% (1-shot), a 3.51% improvement over DCAMA.
This lead is particularly evident in challenging industrial and agricultural domains, where environmental perturbations are severe. On the Road Cracks dataset, our ResNet-50 based AAD achieve 8.15% in the 1-shot setting, while most competing methods, including PFENet and CyCTR, score below 1%. For Steel Defects, AAD (ResNet-101) reaches 15.36% (1-shot), a 5.64% improvement over the strong baseline SCCAN. In the biology and medical domains, AAD also establishes new SOTA results. With a Transformer backbone on the Polyp dataset, AAD achieves a remarkable 32.85% and 51.79% mIoU in 1-shot and 5-shot settings, surpassing the previous best (DCAMA) by 10.53% and 25.51%, respectively. Even in scenarios where other methods perform well, such as the Animal dataset, AAD with ResNet-101 leads with 61.11% mIoU (1-shot), improving upon DCAMA’s 55.51%.
The superior performance of AAD stems from its novel architecture designed to explicitly tackle the environmental variance between support and query images. Unlike traditional FSS methods that rely on direct, and often brittle, feature matching, AAD introduces an adaptive distillation process. By generating abstract class queries and iteratively refining them using foreground information from both support and query images, our model learns to distill the core, invariant semantics of a target category. This allows it to maintain focus on the object’s essential characteristics even in the presence of significant appearance shifts caused by camouflage, motion blur, or occlusion. The Correlation Learner provides a strong initial localization, while the Adaptive Attention Distillation learner purifies this understanding. This leads to a more robust and generalizable representation that excels in the complex, real-world scenarios presented by the ER-FSS benchmark, setting a new standard for robust segmentation.
We present visualization results comparing our method with SOTA approaches in Fig. 5. The qualitative comparisons highlight the limitations of existing methods when faced with the challenging scenarios in the ER-FSS benchmark. For instance, both HMNet and PFENet++ struggle significantly with elongated and small targets, often resulting in a complete failure to locate the target objects in the query image. In contrast, AAD demonstrates superior robustness and adaptability across all these difficult cases. It not only successfully transfers accurate category recognition capabilities to hard samples but also exhibits a remarkable ability to precisely delineate object boundaries. By distilling core, environment-invariant class semantics, AAD effectively ignores background clutter and appearance variations. This allows it to accurately identify and segment challenging targets—whether they are small, elongated, or camouflaged—producing segmentation masks that are both complete and clean, proving its effectiveness in real-world conditions.
@@
Figure 4: Visualization of query maps at different stages with Swin-transformer backbone. After CL refers to the output after the correlation learning module.
@@ V-C Further Analysis
TABLE IV: Ablation Study of correlation learning (CL) module and class discriminative information learner (ADD) with Swin-transformer. The baseline refers to HSNet [32].
@@
ADD 15.36 25.39 34.37 39.10 32.85 51.79
(1) Ablation Study for Proposed Modules. We present experimental results for the proposed elements, the correlation learner (CL) module and ADD. Tab. IV indicates that incorporating CL improves performance by 6%, 17.8%, 14.0%, 18.2%, 19.4%, and 38.5% on three datasets compared to the baseline. The additional inclusion of ADD results in further improvements of 7.6%, 17.8%, 14.0%, 18.2%, 19.4%, and 38.5% over the baseline. It can be proved that ADD enhances the results of both the baseline and baseline+CL, with more pronounced effects as the number of shots increases. This is attributed to ADD’s ability to utilize more support images to generate more accurate class discriminative information.
Moreover, Fig. 4 illustrates the query feature maps after the Backbone, CL, and ADD, demonstrating that the interactive learning of the CL module helps the model disregard background features in certain queries. The addition of ADD significantly enhances the model’s focus on target class information within the query features.
TABLE V: Ablation study for different numbers of random queries with Swin-transformer.
Queries(#) Steel defect Leaf disease Polyp
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
0 7.72 7.56 20.31 20.94 13.48 13.13
5 14.88 19.38 33.53 38.53 36.35 49.31
15 15.36 25.39 34.37 39.10 32.85 51.79
30 14.96 21.40 33.04 37.49 31.80 43.39
50 13.83 20.89 33.13 35.64 36.51 51.26
100 14.82 14.23 32.36 42.55 31.15 32.99
(2) Ablation Study for Numbers of Random Queries. Tab. V demonstrates the impact of varying numbers of learnable queries when the backbone is the Swin-Transformer. From the table, it is observed that as the number of queries increases from $5$, there is minimal difference in the 1-shot results, oscillating within a range of approximately $\pm 2\%$. The 5-shot results show an initial increase followed by a decline. Considering the increased computational resources associated with a higher number of queries, we choose 15 as the base parameter for ADD model.
@@
ADD 32.85 51.79 56.64 59.68 61.09 62.25
(5) Comparison on More Settings of Shots. Also, we provide comparison results between our method and SOTA approaches in various shot settings, as shown in Tab. VIII. It demonstrates that ADD outperforms the second-ranked DCAMA by 10.6%, 25.51%, 27.82%, 27.98%, 27.99%, and 27.30% in 1-shot, 5-shot, 10-shot, 20-shot, 30-shot, and 50-shot settings, respectively. This confirms that ADD effectively utilizes the given support images, and the performance steadily improves as the shot number increases. In contrast, FCN and SAM show relatively minor performance improvements as the shot number increases, indicating that transfer learning-based pre-trained segmentation models struggle to transfer knowledge from simple samples to hard ones effectively.
@@
(6) K-shot Inference. Additionally, when facing $K$-shot Inference, there are two commonly used methods: the voting method, which averages predictions for each support image and query separately, and the average method which averages the features of $K$ support images to generate a segmentation result for the query. Tab. IX displays the results of these two methods on ADD and compares their testing times in the 20-shot scenario. It indicates that the averaging method takes less time and outperforms the other by 6.4% on the Polyp. The differences between the two testing methods are minimal for the other three datasets. In summary, the averaging method provides a higher cost-effectiveness.

### v2

@@ IV-F Decoder and Loss Function
The final stage of the AAD framework employs a decoder to synthesize the coarse localization from the Correlation Learner with the high-level semantic knowledge distilled by the AAD learner. As illustrated in Fig. 3(d), the inputs are the multi-scale coarse query masks $M^{q^{\prime}}_{i}$ and the final set of discriminative class queries $\hat{q}$. To effectively integrate these inputs, we first refine the coarse masks via element-wise multiplication with $\hat{q}$. This operation re-weights the spatial features, amplifying regions consistent with the distilled class semantics while suppressing background noise. The refined features are then concatenated with the original coarse masks to yield a fused representation $R_{q}$:
@@
This fused representation is processed by a decoder composed of three sequential mixing modules. Each module primarily utilizes two convolutional layers coupled with ReLU activations to progressively compress channel dimensions (from $256+\text{num\_queries}\to 128\to 64\to 16\to 2$) while refining spatial details. Through iterative upsampling, the decoder restores the feature maps to the original image resolution to generate the final prediction. The entire framework is trained end-to-end by minimizing the standard Binary Cross-Entropy (BCE) loss between the predicted segmentation map and the ground-truth query mask.
@@ V Experiments
ResNet-50 PFENet  [TPAMI2022] [46] 0.44 0.41 0.39 9.06 9.06 9.14 11.60 11.06 11.13
@@
ResNet-101 DCAMA [ECCV2022] [42] 1.55 1.60 1.62 8.08 9.06 17.52 22.91 26.47 28.23
@@
Matcher [ICLR2024] [26] 1.62 1.89 7.76 8.17 15.23 15.26
MGF-SAM [NIPS2024] [56] 3.66 5.05 4.84 8.99 10.32 22.07 15.81 21.09 23.54
GRPN [AAAI2025] [37] 10.48 10.37 10.64 12.50 18.18 30.85 20.44 27.37 29.10
FCP [AAAI2025] [35] 2.77 2.99 3.08 9.14 13.62 14.61 17.73 17.85 17.81
FS-SAM [ICML2025] [53] 3.94 3.72 11.22 8.31 20.32 20.32 26.21 35.62 38.91
@@
ResNet-50 PFENet  [TPAMI2022] [46] 32.89 33.43 32.74 9.06 9.06 9.14 13.48 13.73 13.81
@@
ResNet-101 DCAMA [ECCV2022] [42] 55.51 59.35 59.06 10.20 10.75 11.21 21.71 25.13 31.85
@@
Matcher [ICLR2024] [26] 9.30 9.32 9.30 9.52 15.00 14.78
MGF-SAM [NIPS2024] [56] 9.44 10.09 11.40 9.13 9.22 9.25 14.96 14.62 14.37
GRPN [AAAI2025] [37] 38.66 43.22 43.89 10.99 11.18 11.16 18.29 18.49 18.62
FCP [AAAI2025] [35] 42.67 45.81 50.47 8.96 8.97 8.97 12.75 12.56 12.50
FS-SAM [ICML2025] [53] 57.09 59.06 69.54 9.25 9.42 9.16 17.76 21.18 21.72
@@
TABLE III: Comparison with SOTA methods on 1-shot, 5-shot, and 20-shot setting on Lunar terrain (Astronomy), and City satellite (Geography) datasets. The ’Mean’ refers to the average results across all eight datasets in the ER-FSS benchmark. The numbers in Bold indicate the best performance.
@@
ResNet-50 PFENet  [TPAMI2022] [46] 7.02 7.65 7.98 4.91 4.66 4.50 11.06 11.13 11.10
@@
ResNet-101 DCAMA [ECCV2022] [42] 4.91 7.62 8.23 9.48 10.00 11.25 16.79 20.03 21.12
@@
Matcher [ICLR2024] [26] 7.76 8.17 3.66 3.86 8.93 9.05
MGF-SAM [NIPS2024] [56] 8.99 10.32 12.07 4.41 5.69 6.47 9.84 11.29 12.07
GRPN [AAAI2025] [37] 13.92 16.59 18.07 7.66 8.16 8.04 16.61 19.19 21.29
FCP [AAAI2025] [35] 3.39 3.37 3.25 8.86 8.87 8.85 13.28 12.92 14.94
FS-SAM [ICML2025] [53] 11.58 15.16 18.87 9.20 11.55 15.28 17.92 22.00 25.63
@@ V-A Experiment Setup
Training and Testing Strategy. We evaluate two categories of methods using distinct protocols tailored to their architectural paradigms. For pre-trained segmentation models (e.g., SAM [20], Matcher [26], MGF-SAM [56], FS-SAM [53]), we follow a transfer learning strategy: the models undergo end-to-end pre-training on the base dataset, and during evaluation, they are fine-tuned on the support images before predicting query masks. For FSS-specific models (e.g., PFENet [46], DCAMA [42], HDMNet [36], SCCAN [52], PFENet++ [31], HMNet [51], ABCDFSS [18], NTRENet++ [27], GRPN [37]), we adopt the standard meta-learning paradigm for both training and testing. Evaluation follows the protocol in [24]: we report the average mean-IoU over 2 runs [47] with different random seeds, where each run consists of 1,000 tasks per dataset to ensure statistical reliability.
@@
Implementation Details. For a fair comparison, we employ Swin-transformer [28], ResNet-50 [17], and ResNet-101 [17] as feature extraction networks, all of which are initialized with the weights pre-trained on ILSVRC [39] and kept frozen during the training process, following the previous works [24, 2]. The input dimensions for support and query images are set at $384\times 384$. ResNet-50 and ResNet-101 feature maps have channel dimensions of 256, 512, 1024, and 2048, while Swin-transformer feature maps have dimensions of 192, 384, 768, and 1536. For ResNet-50-based AAD, the number of learnable initialization queries is set to 15, and the interaction with $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ dimensions of support and query features. In the case of the ResNet-101-based model, the number of queries is set to 20, with feature interactions at $\frac{1}{32}$ dimension Meanwhile, for the Swin-transformer-based model, 15 queries are used, with feature interactions at $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ dimensions. The decoder is configured with $2$ convolutional layers, and between each module, bilinear interpolation is applied to upsample the feature maps by a factor of $2$, resulting in a total of $2$ upsampling functions. These networks were implemented using PyTorch, with AdamW [30] as the optimizer, a learning rate of 1e41e-4, and a weight decay of $0.05$. During training, the batch size is set to $120$, and the training process ran on $8$ NVIDIA A800-SXM4-80GB GPUs in parallel, with subsequent evaluation on one GPU.
@@ V-B Comparison with SOTA Models
We present a comprehensive comparison of our proposed AAD framework against SOTA methods across eight diverse datasets in the ER-FSS benchmark. As summarized in Table III (Mean), AAD consistently outperforms existing approaches across all backbone architectures and shot settings. Notably, with the ResNet-50 backbone, AAD achieves average mean-IoU scores of 19.24%, 21.90%, and 23.84% for 1-shot, 5-shot, and 20-shot tasks, respectively. This represents a substantial improvement over the previous best-performing method, NTRENet++ [27] (e.g., +10.24% in the 1-shot setting). Similar trends are observed with the ResNet-101 backbone, where AAD surpasses the strong baseline DCAMA [42] by margins of 4.43%, 5.00%, and 6.09% across the respective shot settings. These results validate the efficacy of our adaptive attention mechanism in learning generalized representations.
Furthermore, AAD demonstrates superior robustness in challenging environments characterized by high noise and domain shifts, such as industrial defects and medical imaging. As shown in Table I, on the Steel defect dataset—which features subtle targets and complex textures—AAD with ResNet-50 achieves a 1-shot IoU of 10.44%, significantly outperforming NTRENet++ (4.41%) and ABCDFSS (7.90%). Similarly, in the medical domain (Table II), AAD excels on the Polyp dataset, achieving a remarkable 21.93% in the 1-shot setting compared to 18.29% for NTRENet++. These gains highlight the model’s ability to suppress environmental noise and accurately localize targets even when visual cues are obscure, effectively mitigating the domain gap between training and testing distributions.
Finally, we compare AAD against recent Transformer-based and SAM-based methods, which typically benefit from large-scale pre-training. Despite the strong transfer capabilities of models like SAM [20] and FS-SAM [53], AAD maintains a competitive edge. On the Animal dataset (Table II), AAD achieves a 1-shot performance of 63.26%, surpassing FS-SAM (57.09%) and GRPN (38.66%). Even in the aggregate mean results (Table III), our Transformer-based variant leads the leaderboard with 23.47% (1-shot) and 29.62% (5-shot), outperforming the sophisticated FS-SAM by 5.55% and 7.62%, respectively. This superiority suggests that while pre-training provides a strong foundation, our proposed attention distillation mechanism is crucial for adapting to specific, unseen tasks where intra-class variance and environmental interference are prevalent.
@@
Figure 4: Visualization of query maps at different stages under various environmental perturbations using the Swin-Transformer backbone. “After CL” denotes the output following the correlation learning module.
@@ V-C Further Analysis
(1) Ablation Study for Proposed Modules. We present experimental results for the proposed elements, the correlation learner (CL) module and AAD. Tab. IV indicates that incorporating CL improves performance by 6%, 17.8%, 14.0%, 18.2%, 19.4%, and 38.5% on three datasets compared to the baseline. The additional inclusion of AAD results in further improvements of 7.6%, 17.8%, 14.0%, 18.2%, 19.4%, and 38.5% over the baseline. It can be proved that ADD enhances the results of both the baseline and baseline+CL, with more pronounced effects as the number of shots increases. This is attributed to AAD’s ability to utilize more support images to generate more accurate class discriminative information.
TABLE IV: Ablation Study of correlation learning (CL) module and class discriminative information learner (AAD) with Swin-transformer. The baseline refers to HSNet [33].
@@
AAD 15.36 25.39 34.37 39.10 32.85 51.79
Moreover, Fig. 4 visualizes the evolution of query feature maps at three critical stages: post-Backbone, post-Correlation Learner (CL), and post-AAD. These visualizations cover a spectrum of challenging environmental perturbations, including image blur, small and slender objects, occlusion, and camouflage. Observing the progression, it is evident that while the interactive learning within the CL module offers a preliminary filtration of irrelevant background noise compared to the raw backbone features, it often struggles to fully isolate the target in complex scenarios. In contrast, the integration of the AAD module proves decisive. By distilling robust class semantics, AAD significantly sharpens the model’s focus, enabling it to accurately localize and highlight difficult targets even when they are obscured by camouflage or severe occlusion, thereby validating its superior environmental robustness.
TABLE V: Ablation study on the effectiveness and efficiency of the number of learnable queries.
Queries Steel defect Leaf disease Polyp GFLOPs
num(#) 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
0 7.72 7.56 20.31 20.94 13.48 13.13 275
5 14.88 19.38 33.53 38.53 36.35 49.31 275
15 15.36 25.39 34.37 39.10 32.85 51.79 275
30 14.96 21.40 33.04 37.49 31.80 43.39 276
50 13.83 20.89 33.13 35.64 36.51 51.26 276
100 14.82 14.23 32.36 42.55 31.15 32.99 277
(2)The Effectiveness and Efficiency of the Number of Learnable Queries. Table V investigates the impact of the number of learnable queries ($N$) using the Swin-Transformer backbone. Performance remains relatively stable across variations: 1-shot results fluctuate within a narrow $\pm 2\%$ range, while 5-shot results exhibit a slight rise followed by a decline. Consequently, we select $N=15$ as the optimal setting. Importantly, increasing $N$ (up to 100) incurs negligible computational overhead. Since the queries function as low-dimensional vectors interacting within the lightweight AAD module, their contribution to the total GFLOPs and inference time is minimal compared to the backbone feature extraction. Thus, the model achieves robust performance with high efficiency regardless of minor variations in query count.
@@
AAD 32.85 51.79 56.64 59.68 61.09 62.25
(5) Comparison on More Settings of Shots. Also, we provide comparison results between our method and SOTA approaches in various shot settings, as shown in Tab. VIII. It demonstrates that AAD outperforms the second-ranked DCAMA by 10.6%, 25.51%, 27.82%, 27.98%, 27.99%, and 27.30% in 1-shot, 5-shot, 10-shot, 20-shot, 30-shot, and 50-shot settings, respectively. This confirms that AAD effectively utilizes the given support images, and the performance steadily improves as the shot number increases. In contrast, FCN and SAM show relatively minor performance improvements as the shot number increases, indicating that transfer learning-based pre-trained segmentation models struggle to transfer knowledge from simple samples to hard ones effectively.
@@
(6) K-shot Inference. Additionally, when facing $K$-shot Inference, there are two commonly used methods: the voting method, which averages predictions for each support image and query separately, and the average method which averages the features of $K$ support images to generate a segmentation result for the query. Tab. IX displays the results of these two methods on AAD and compares their testing times in the 20-shot scenario. It indicates that the averaging method takes less time and outperforms the other by 6.4% on the Polyp. The differences between the two testing methods are minimal for the other three datasets. In summary, the averaging method provides a higher cost-effectiveness.